OpenStudy (anonymous):

HELP!! PLEASE! f(x) = quantity x squared plus x minus two divided by quantity x squared minus three x minus four Graph A coordinate axis scaled by one. Domain and Range (1 point): _____________________ x and y Intercept(s) (1 point): _____________________ Horizontal Asymptote(s) (1 point): ___________________ Vertical Asymptote(s) (1 point): ____________________

OpenStudy (solomonzelman):

\(\Large\color{black}{ f(x)=\frac{x^2+x-2}{x^2-3x+4} }\) . right?

OpenStudy (solomonzelman):

first find any domain restrictions (when x is equal to zero) and that will be our vertical asymptote(s).

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Large\color{black}{ x^2-3x+4\ne0 }\) .

OpenStudy (anonymous):

Im still I little confused, give me a minute.

OpenStudy (anonymous):

Ok, so would our Vertical asymptote: 0?

OpenStudy (anonymous):

im lost

OpenStudy (solomonzelman):

can you solve for x, when x^2-3x+4=0?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

x^2-3x+4=0 (x-3) (x+4)

OpenStudy (anonymous):

just got off a concussion sorry.

OpenStudy (solomonzelman):

you can see that x^2-3x +4 can't be zero, because x^2-3x, for any smallest value of x, would have a lesser absolute value than 4.

OpenStudy (solomonzelman):

So you don't really have an asymptote of a real number. I mean a vertical asymptote.

OpenStudy (solomonzelman):

Okay, knowing that you don't have a vert. asymptote...

OpenStudy (anonymous):

yea, so its 0 right

OpenStudy (solomonzelman):

close lol. I would say, there is ZERo vertical asymptotes. Look over what I said, if you don't know why is that...

OpenStudy (solomonzelman):

Tell me when you are good to proceed.

OpenStudy (anonymous):

I just said, 0 vertical asymptotes, but ok proceed.

OpenStudy (solomonzelman):

Do you know why there is no vertical asymptote? are you sure you want to go on?

OpenStudy (anonymous):

yes, please go on.

OpenStudy (solomonzelman):

Okay, so the domain, since there is no vertical asymptote - which removes any possible domain restriction, is all real numbers (you know what that means in an interval notation, right?) The range however, we have to solve for. \(\large\color{black}{ y=\frac{\LARGE x^2+x-2}{\LARGE x^2-3x+4} }\) \(\large\color{brown}{ (x^2-3x+4)y= x^2+x-2 }\)

OpenStudy (solomonzelman):

do you see just the math that I did so far?

OpenStudy (solomonzelman):

I multiplied both sides times `x^2-3x+4` right?

OpenStudy (anonymous):

oh ok.

OpenStudy (solomonzelman):

I lost my latex work...

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/lq6z91wssn TAKE A LOOK AT THIS GRAPH IF YOU WANT...

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

ther eis your range, it goes from 0.5 to 2.094

OpenStudy (solomonzelman):

No x intercepts, because the range never reaches zero.

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

actually there is a bit lower point where range is 0.478 (a but lower than 0.5) but the y-intercept is (0, 0.5)

OpenStudy (solomonzelman):

there is no horizontal asymptote either.

OpenStudy (anonymous):

would the horizontal asymptote be 0 or is there none at all?

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

yes, none at all.

OpenStudy (anonymous):

So what is the domain again